OpenStudy (anonymous):

simplify this? (2n^4h^3)^4

OpenStudy (shamil98):

The formula or rule for this is \[\huge (a^m)^n = a^{m*n}\]

OpenStudy (mathstudent55):

Also, \( (abc)^n = a^nb^nc^n \)

OpenStudy (anonymous):

So it'd be \[2n ^{4}h ^{3*4}\] ?

OpenStudy (mathstudent55):

Every factor inside the parentheses, including the 4, is raised to teh 4th power.

OpenStudy (shamil98):

\[\huge 2^4 * n^{4*4} * h^{3*4}\] My bad.

OpenStudy (anonymous):

okay, so it would be? \[2^{4}n ^{16}h ^{12}\]

OpenStudy (shamil98):

Yes, but you can simplify 2^4 further.

OpenStudy (anonymous):

2^4 equals 8, so now is that it simplifies to?

OpenStudy (mathstudent55):

You are correct. \((2n^4h^3)^4\) Use the second rule first. When a product is raised to a power, raise every factor to that power. \(= 2^4 \cdot (n^4)^4 \cdot (h^3)^4 \) Now use the first rule. \(=16n^{16}h^{12}\)

OpenStudy (anonymous):

Thanks so much :D

OpenStudy (mathstudent55):

wlcm

OpenStudy (shamil98):

Just to clarify. 2^4 = 16 2 x 2 x 2 x 2 = 16 not 8.

OpenStudy (anonymous):

OH lol right sorry I was thinking 2^3 xp THANKS!